OpenStudy (anonymous):

Find equations of the lines that pass through the point (1,4) and have the given slope: a. 1 b. -1 c. 3 d. 1/3 Can someone just please explain this question..?

OpenStudy (anonymous):

Use the equation y-y1=m(x-x1) For example: a)y-1=1(x-4) m=slope y1=given y on your point x1=given x on your point

OpenStudy (anonymous):

oh ok yea. I thought so. Thanks a lot!